A spring extends 2 inches when loaded vertically with 4 lbs. This weight was pulled further down by 3 inches an was released from rest. Determine the frequency of vibration in Hz


1
Expert's answer
2021-04-01T01:12:43-0400

Given as,

Mass attached (m)=4lbs

Extension in the spring(y)=3inches=0.25 ft

The spring constant(k)="\frac{m×g}{y}=\frac{4×32.2}{0.25}=515.2\: lb-f/ft"

The natural frequency is given by,

"f= \frac{1}{2\pi} \sqrt{\frac{k}{m}}=1.81 Hz"